# Stochastic resonance in vibrational polariton chemistry

### v1

@@ I Introduction
Naturally, this intriguing yet controversial phenomenon has spurred a substantial amount of theoretical research.Galego, Garcia-Vidal, and Feist (2016, 2017); Galego et al. (2019); Campos-Gonzalez-Angulo and Yuen-Zhou (2020); Mandal, Montillo Vega, and Huo (2020); Yang and Cao (2021); Li, Mandal, and Huo (2021a); Li, Nitzan, and Subotnik (2020); Li, Mandal, and Huo (2021b); Sun and Vendrell (2022); Schäfer et al. (2022); Mandal, Li, and Huo (2022); Lindoy, Mandal, and Reichman (2022); Wang et al. (2022); Fischer, Anders, and Saalfrank (2022); Fiechter et al. (2023); Sokolovskii and Groenhof (2023); Pavošević, Smith, and Rubio (2023) Recent studies using quantum dynamical simulations underscore the necessity of a quantum discrete state description for both molecular vibrations and the cavity mode.Lindoy, Mandal, and Reichman (2023, 2024); Ying and Huo (2023); Hu, Ying, and Huo (2023); Ke, Borrelli, and Thoss (2022); Ke and Richardson (2024) However, this requirement introduces significant challenges, particularly in the collective regime, where a large ensemble of molecules couples to the cavity mode. To address the computational challenges, we employ the hierarchical equations (HEOM) of motion along with a tree tensor network state (TTNS) solver.Ke (2023) The HEOM method is a well-established, numerically exact approach for open quantum system dynamics,Tanimura and Kubo (1989); Yan et al. (2004); Ishizaki and Tanimura (2005); Xu and Yan (2007); Shi et al. (2009); Yan (2014); Jin, Zheng, and Yan (2008); Schinabeck, Härtle, and Thoss (2018); Hsieh and Cao (2018); Shi et al. (2018); Tanimura (2020) enabling a non-perturbative and non-Markovian treatment of system dynamics and bath-related observables. The TTNS,Shi, Duan, and Vidal (2006); Tagliacozzo, Evenbly, and Vidal (2009); Murg et al. (2010); Li, von Delft, and Xiang (2012); Changlani et al. (2013); Nakatani and Chan (2013); Lubich et al. (2013); Murg et al. (2015); Gunst et al. (2018); Schröder et al. (2019); Larsson (2019); Ferrari, Magnifico, and Montangero (2022); Montangero, Rico, and Silvi (2022); Sulz et al. (2024) in addition, provides an efficient data compression scheme for storing and propagating the high-dimensional composite system-plus-bath wavefunction.
@@
To gain a preliminary understanding of how chemical reactions inside a cavity respond to ambient noise–stemming from interactions between the confined cavity mode and the continuum of far-field electromagnetic modes, which dampens the cavity’s oscillatory dynamics–we begin our study on the single-molecule level. Our results show that reaction rate enhancements within the cavity exhibit a typical stochastic resonance featureGammaitoni et al. (1998): neither excessively weak nor overly strong damping of cavity dynamics supports effective cavity-induced rate modifications. Instead, the rate constant inside a resonant cavity reaches its peak at an intermediate level of cavity damping. In the collective coupling regime, where multiple molecules are interconnected through a shared coupling with the cavity mode, rate modifications can display distinct behaviors as the system size grows. Under weak ambient noise, the rate enhancement strengthens as more molecules are coupled to the cavity mode. On the contrary, under conditions of strong cavity damping, the rate enhancement is attenuated with the increased number of coupled molecules. These observations lead us to ask an essential question: Could an ensemble of molecules on the macroscopic scale, immersed in a solvent and thus exposed to significant background noise, act cooperatively through collective coupling to the cavity mode, and synergetically harness a feeble external fluctuation to optimize the reaction efficiency?
@@ II Theory
where the cavity mode and each molecule are coupled to their respective baths, represented as collections of harmonic oscillators. The bath associated with the cavity mode is referred to as the cavity bath, and each molecular bath is referred to as a solvent bath. The $k$th oscillator in bath $\alpha$ (either cavity or solvent) is described by its momentum $P_{\alpha k}$ and coordinate $Q_{\alpha k}$, with frequency $\omega_{\alpha k}$ and coupling strength $g_{\alpha k}$. The coupling to the system induces a displacement $g_{\alpha k}x_{\alpha}/\omega_{\alpha k}^{2}$ in each oscillator’s coordinate.
@@
$C_{\alpha}(t)=\frac{1}{\pi}\int_{-\infty}^{\infty}\frac{e^{-i\omega t}}{1-e^{-\beta\omega}}J_{\alpha}(\omega)\mathrm{d}\omega=\sum_{p=1}^{P\rightarrow\infty}\lambda_{\alpha}\eta_{\alpha p}e^{-i\gamma_{\alpha p}t},$ (4)
where $\beta=1/k_{\rm B}T$ is the inverse temperature, and $J_{\alpha}(\omega)=\frac{\pi}{2}\sum_{k}\frac{c_{\alpha k}^{2}}{\omega_{\alpha k}}\delta(\omega-\omega_{\alpha k})$ is the spectral density function of bath $\alpha$. This spectral density encodes the coupling-weighted density of states in the frequency domain for each bath. The time correlation function $C_{\alpha}(t)$ can be expanded analytically or numerically as a series of exponentials.Hu, Xu, and Yan (2010); Xu et al. (2022) This exponential expansion can be understood as mapping the original bath onto an effective one, which can be truncated efficiently at finite temperatures. In this effective representation, the bath consists of $P$ dissipative bosonic modes, each with a complex-valued frequency $\gamma_{\alpha p}$ and a component-wise coupling strength $\eta_{\alpha p}$ to the system. The reorganization energy $\lambda_{\alpha}=\sum_{k}\frac{g_{\alpha k}^{2}}{2\omega_{\alpha k}^{2}}$ quantifies the overall coupling strength between the system and bath $\alpha$.
@@
$\begin{split}\mathcal{H}=&\hat{H}_{\rm S}+\sum_{\alpha}\lambda_{\alpha}\hat{x}_{\alpha}^{2}-\tilde{H}_{\rm S}-\sum_{\alpha}\lambda_{\alpha}\tilde{x}_{\alpha}^{2}-i\sum_{\alpha}\sum_{p}\gamma_{\alpha p}b_{\alpha p}^{+}b_{\alpha p}\\ &+\sum_{\alpha}\sum_{p}\sqrt{\lambda_{\alpha}}\left[(\hat{x}_{\alpha}-\tilde{x}_{\alpha})b_{\alpha p}+(\eta_{\alpha p}\hat{x}_{\alpha}-\eta^{*}_{\alpha p}\tilde{x}_{\alpha})b_{\alpha p}^{+}\right].\end{split}$ (8)
@@
To facilitate efficient propagation of the Schrödinger equation in Eq. (5), we employ a tree tensor network state solver.Ke (2023) The high-rank coefficient tensor $C^{\mathbf{n}}_{v_{\rm c}v^{\prime}_{\rm c},\cdots v_{N_{\rm mol}}v^{\prime}_{N_{\rm mol}}}(t)$ is decomposed as a tree tensor network state, which is a powerful way to handle the complicated tensor structure involved in this system. As an example, a graphical representation of the TTNS decomposition of $|\Psi(t)$ for a system with $N_{\rm mol}=8$ molecules and $P=4$ effective environmental modes per bath is shown in Fig. 1(b). In this decomposition, a node with $z$-legs corresponds to a rank-$z$ tensor. To achieve an optimal balance between numerical stability, computational efficiency, and simulation time, the number of legs $z$ is constrained to $z\leq 3$. The physical indices $v_{\rm c}/v^{\prime}_{\rm c}$ (for the cavity), $v_{i}/v^{\prime}_{i}$ (for the molecules), and $n_{{\rm c}p},\,n_{ip}$ (for the effective environmental modes) in $|\Psi(t)\rangle$ are represented by the dangling legs in the yellow (filled/hollow), red (filled/hollow), blue, and green nodes, respectively. The connected legs between neighboring nodes $j$ and $k$ correspond to virtual indices $r_{jk}$, which runs from $1$ to $D_{jk}$. The maximum value among the bond dimensions $\{D_{jk}\}$ is called the maximal bond dimension $D_{\rm max}$. In practice, $D_{\rm max}$ is systematically increased until the results converge to a desired accuracy. The gray nodes, which have exclusively connected legs, are termed connecting nodes. The root node is assigned to the filled yellow node for the cavity, which has no parent node. All other nodes, except for the root node, have a parent node. This parent node is the adjacent node on the path pointing toward the root. In the binary tree structure displayed in Fig. 1(b), each node has at most two children. The leaves are the nodes without any offspring, representing the termination points of the tree.
@@
The super-Hamiltonian $\mathcal{H}$ can be represented as a tree tensor network operator (TTNO) with the same tree topology as the TTNS used for $|\Psi(t)\rangle$. This decomposition ensures that when the TTNO is applied to the TTNS of $\Psi(t)$, the tree structure is preserved. The construction of the TTNO for $\mathcal{H}$ can be performed in an automatic manner.Milbradt, Huang, and Mendl (2024a, b); Ceruti, Kressner, and Sulz (2024); Li et al. (2024) In this work, we use an optimal TTNO constructor based on bipartite theory, which allows for an efficient construction of the TTNO.Li et al. (2024) The resulting TTNO, corresponding to the tree shape shown in Fig. 1(b), is depicted in Fig. 1(d). A key difference between the TTNS and the TTNO is that in the TTNO, the physical nodes possess two dangling legs. To evolve Eq. (5), we adopt a time propagation scheme based on the time-dependent variational principle (TDVP), as detailed in Refs. Haegeman et al., 2016; Paeckel et al., 2019; Dunnett and Chin, 2021; Bauernfeind and Aichhorn, 2020; Kloss, Reichman, and Bar Lev, 2020; Ceruti, Lubich, and Walach, 2021. After solving Eq. (5), observables of interest can be extracted from the wavefunction $|\Psi(t)\rangle$.
@@
To perform the inner product in Eq. (9), the wavefunction $|{\rm II}\rangle$ is also decomposed as a TTNS, maintaining the same tree topology as $|\Psi(t)\rangle$, as displayed in Fig. 1(c). The number next to the legs denotes the bond dimension of the corresponding index. The parameters $d_{\rm m}$, $d_{\rm c}$, and $d_{\rm e}$ represent the number of basis functions for the molecular reactive DoF, cavity mode, and the effective environmental bosons, respectively. Each node in the TTNS is represented by a rank-$3$ tensor $V_{i_{1}\times i_{2}\times i_{3}}$, where $i_{1}$, $i_{2}$, and $i_{3}$ are the dimensions of the respective indices. For the connecting nodes, the tensor is $V_{1\times 1\times 1}=1$. For the molecular (in red) or cavity (in yellow) nodes, the non-zero entries in the tensor are $V_{1\times d_{\rm m/c}\times d_{\rm m/c}}[1,j,k]=\delta_{jk}$ (filled node) and $V_{d_{\rm m/c}\times 1\times d_{\rm m/c}}[j,1,k]=\delta_{jk}$ (hollow node). For the environmental nodes (in green and blue), the non-zero entry is $V_{1\times 1\times d_{\rm e}}[1,1,1]=1$.
@@
In the simulations conducted for this study, the following parameters are used for the double-well model, $E_{b}=2250\,\text{cm}^{-1}$ and $a=44.4$ a.u. , which are consistent with previous studies in the field.Lindoy, Mandal, and Reichman (2023, 2024); Ying and Huo (2023); Hu, Ying, and Huo (2023); Ke, Borrelli, and Thoss (2022); Ke and Richardson (2024); Fiechter et al. (2023) The Debye-Lorentzian spectral density function is adopted to characterize the baths and their interaction with the system
$J_{\alpha}(\omega)=\frac{2\lambda_{\alpha}\omega\Omega_{\alpha}}{\omega^{2}+\Omega_{\alpha}^{2}},$ (13)
@@
Figure 2: Ratio of rates $k_{\rm ic}/k_{\rm oc}$ for a single molecule undergoing a reaction inside an optical cavity $(k_{\rm ic})$ compared to the outside rate $(k_{\rm oc})$ as a function of the cavity damping strength $\lambda_{\rm c}$. The results for two different temperatures $T$, characteristic frequencies $\Omega_{\rm c}$ of the cavity bath, and light-matter coupling strengths ($\eta_{\rm c}=0.00125$ a.u. in the upper panel and $\eta_{\rm c}=0.005$ a.u. in the lower panel) are shown. The cavity frequency is kept at $\omega_{\rm c}=1185\,\text{cm}^{-1}$ and $d_{\rm c}=10$.
@@ III Results
In practice, the cavity is not perfectly reflective and is therefore subject to external noise from the continuum of far-field electromagnetic modes. Previous theoretical studies have demonstrated that, in a resonant but lossless cavity, reaction rates remain unaffected.Lindoy, Mandal, and Reichman (2023); Ying and Huo (2023); Ke, Borrelli, and Thoss (2022) This finding gives a hint that cavity leakage may be a crucial factor in altering reaction dynamics inside an optical microcavity. Accordingly, the objective of this work is to understand how ground-state chemical reactions within an optical cavity are influenced by cavity damping. We begin by investigating the effect of cavity damping strength on reaction rate modifications at the single-molecule level, then extend our analysis to the collective regime, where multiple molecules are simultaneously coupled to the cavity mode.
In this work, cavity leakage is modeled by introducing a cavity bath, with the cavity damping strength quantified by the parameter $\lambda_{\rm c}$. This is verified by examining the dynamics of the average photonic excitation number, given by $\langle n_{\rm c}(t)\rangle=\mathrm{Tr}\{a^{\dagger}_{\rm c}a_{\rm c}\rho(t)\}$. Here, $a^{\dagger}_{\rm c}=\sqrt{\omega_{\rm c}/2}(x_{\rm c}-ip_{\rm c/\omega_{\rm c}})$ and $a_{\rm c}=\sqrt{\omega_{\rm c}/2}(x_{\rm c}+ip_{\rm c/\omega_{\rm c}})$ represent the creation and annihilation operators of the confined cavity mode, respectively.
@@
Figure 4: Same as Fig. 3 expect that $\lambda_{\rm c}=165\,\text{cm}^{-1}$.
@@
Fig. 2 presents the cavity-induced rate modification, represented as $k_{\rm ic}/k_{\rm oc}$–the ratio of reaction rates inside and outside the cavity–in the single-molecule limit, plotted as a function of $\lambda_{\rm c}$. The cavity frequency is set at $\omega_{\rm c}=1185\,\text{cm}^{-1}$, closely resonant with molecular vibrational transitions and corresponding to the peak in the rate modification profile over $\omega_{\rm c}$ (see also Fig. 3b and Fig. 4b). The rate within the cavity initially increases with $\lambda_{\rm c}$, reaching a maximum before this rate enhancement diminishes as $\lambda_{\rm c}$ continues to increase. This behavior is examined with various values of temperature $T$, the characteristic frequency $\Omega_{\rm c}$ of the cavity bath, and the light-matter coupling strength $\eta_{\rm c}$. The extent of the rate modification depends on temperature, $\Omega_{\rm c}$, and $\eta_{\rm c}$. Notably, the cavity-induced rate modification is a finite-temperature effect, vanishing at low temperatures and saturating as temperature rises, as demonstrated in our previous study.Ke, Borrelli, and Thoss (2022) Furthermore, the rate modification increases with $\Omega_{\rm c}$ until it reaches a saturation point. However, both temperature $T$ and $\Omega_{\rm c}$ only slightly affect the crossover value of $\lambda_{\rm c}$. In contrast, increasing the light-matter coupling strength $\eta_{\rm c}$ not only strengthens the rate modification significantly but also shifts the crossover $\lambda_{\rm c}$ to a much higher value, rendering the rate modification more resilient to ambient noise.
@@
This observation fits into the notion of the stochastic resonance theory, which was first proposed to explain the periodic recurrence of Earth’s ice ages.Benzi, Sutera, and Vulpiani (1981); Nicolis (1981) Over time, it has prospered into a broad field on its own with diverse applications across physics, chemistry, biomedical sciences, engineering, and beyond, highlighting the constructive role of an optimal amount of noise.Wiesenfeld and Moss (1995); Gammaitoni et al. (1998); Andò and Graziani (2000); Hänggi (2002); Harmer, Davis, and Abbott (2002); Wellens, Shatokhin, and Buchleitner (2003); Moss, Ward, and Sannita (2004) Stochastic resonance refers to a phenomenon where a moderate level of noise improves a system’s sensitivity to detect weak, information-carrying signals.Gammaitoni et al. (1998) For chemical reactions inside an optical cavity, three key elements have been possessed for the emergence of stochastic resonance. Firstly, reactions take place in a solvent, which serves as a source of substantial background noise. Secondly, reactive molecules are inherently nonlinear systems that support threshold-like barrier-crossing events. Thirdly, cavity leakage exposes reactions to a generically weak coherent input, which could boost the system’s sensitivity.

### v2

@@ I Introduction
Naturally, this intriguing yet controversial phenomenon has spurred a substantial amount of theoretical research.Galego, Garcia-Vidal, and Feist (2016, 2017); Galego et al. (2019); Campos-Gonzalez-Angulo and Yuen-Zhou (2020); Mandal, Montillo Vega, and Huo (2020); Yang and Cao (2021); Li, Mandal, and Huo (2021a); Li, Nitzan, and Subotnik (2020); Li, Mandal, and Huo (2021b); Sun and Vendrell (2022); Schäfer et al. (2022); Mandal, Li, and Huo (2022); Lindoy, Mandal, and Reichman (2022); Wang et al. (2022); Fischer, Anders, and Saalfrank (2022); Fiechter et al. (2023); Sokolovskii and Groenhof (2024); Pavošević, Smith, and Rubio (2023) Recent studies using quantum dynamical simulations underscore the necessity of a quantum discrete state description for both molecular vibrations and the cavity mode.Lindoy, Mandal, and Reichman (2023, 2024); Ying and Huo (2023); Hu, Ying, and Huo (2023); Ke and Richardson (2024a, b) However, this requirement introduces significant challenges, particularly in the collective regime, where a large ensemble of molecules couples to the cavity mode. To address the computational challenges, we employ the hierarchical equations (HEOM) of motion along with a tree tensor network state (TTNS) solver.Ke (2023) The HEOM method is a well-established, numerically exact approach for open quantum system dynamics,Tanimura and Kubo (1989); Yan et al. (2004); Ishizaki and Tanimura (2005); Xu and Yan (2007); Shi et al. (2009); Yan (2014); Jin, Zheng, and Yan (2008); Schinabeck, Härtle, and Thoss (2018); Hsieh and Cao (2018); Shi et al. (2018); Tanimura (2020) enabling a non-perturbative and non-Markovian treatment of system dynamics and bath-related observables. The TTNS,Shi, Duan, and Vidal (2006); Tagliacozzo, Evenbly, and Vidal (2009); Murg et al. (2010); Li, von Delft, and Xiang (2012); Changlani et al. (2013); Nakatani and Chan (2013); Lubich et al. (2013); Murg et al. (2015); Gunst et al. (2018); Schröder et al. (2019); Larsson (2019); Ferrari, Magnifico, and Montangero (2022); Montangero, Rico, and Silvi (2022); Sulz et al. (2024) in addition, provides an efficient data compression scheme for storing and propagating the high-dimensional composite system-plus-bath wavefunction.
@@
To gain a preliminary understanding of how chemical reactions inside a cavity respond to ambient noise–stemming from interactions between the confined cavity mode and the continuum of far-field electromagnetic modes, which dampens the cavity’s oscillatory dynamics–we begin our study on the single-molecule level. Our results show that reaction rate enhancements within the cavity exhibit a typical stochastic resonance featureGammaitoni et al. (1998): neither excessively weak nor overly strong damping of cavity dynamics supports effective cavity-induced rate modifications. Instead, the rate constant inside a resonant cavity reaches its peak at an intermediate level of cavity damping. In the collective coupling regime, where multiple molecules are interconnected through a shared coupling with the cavity mode, rate modifications can display distinct behaviors as the system size grows. Under weak ambient noise, the rate enhancement strengthens as more molecules are coupled to the cavity mode. On the contrary, under conditions of strong cavity damping, the rate enhancement is attenuated with the increased number of coupled molecules. These observations lead us to ask an essential question: Could an ensemble of molecules on the macroscopic scale, immersed in a solvent and thus exposed to significant background noise, act cooperatively through collective coupling to the cavity mode, and synergetically harness a feeble fluctuation to optimize the reaction efficiency?
@@ II Theory
where the cavity mode and each molecule are coupled to their respective baths, represented as collections of harmonic oscillators, as schematically illustrated in Fig. 1a. The bath associated with the cavity mode is referred to as the cavity bath, and each molecular bath is referred to as a solvent bath. The $k$th oscillator in bath $\alpha$ (either cavity or solvent) is described by its momentum $P_{\alpha k}$ and coordinate $Q_{\alpha k}$, with frequency $\omega_{\alpha k}$ and coupling strength $g_{\alpha k}$. The coupling to the system induces a displacement $g_{\alpha k}x_{\alpha}/\omega_{\alpha k}^{2}$ in each oscillator’s coordinate.
@@
$C_{\alpha}(t)=\frac{1}{\pi}\int_{-\infty}^{\infty}\frac{e^{-i\omega t}}{1-e^{-\beta\omega}}J_{\alpha}(\omega)\mathrm{d}\omega=\sum_{p=1}^{P\rightarrow\infty}\lambda_{\alpha}^{2}\eta_{\alpha p}e^{-i\gamma_{\alpha p}t},$ (4)
where $\beta=1/k_{\rm B}T$ is the inverse temperature, and $J_{\alpha}(\omega)=\frac{\pi}{2}\sum_{k}\frac{c_{\alpha k}^{2}}{\omega_{\alpha k}}\delta(\omega-\omega_{\alpha k})$ is the spectral density function of bath $\alpha$. This spectral density encodes the coupling-weighted density of states in the frequency domain for each bath. The time correlation function $C_{\alpha}(t)$ can be expanded analytically or numerically as a series of exponentials.Hu, Xu, and Yan (2010); Xu et al. (2022) This exponential expansion can be understood as mapping the original bath onto an effective one, which can be truncated efficiently at finite temperatures. In this effective representation, the bath consists of $P$ dissipative bosonic modes, each with a complex-valued frequency $\gamma_{\alpha p}$ and a component-wise coupling strength $\eta_{\alpha p}$ to the system. The reorganization energy $\lambda_{\alpha}^{2}=\sum_{k}\frac{g_{\alpha k}^{2}}{2\omega_{\alpha k}^{2}}$ quantifies the overall coupling strength between the system and bath $\alpha$.
@@
$\begin{split}\mathcal{H}=&\hat{H}_{\rm S}+\sum_{\alpha}\lambda_{\alpha}^{2}\hat{x}_{\alpha}^{2}-\tilde{H}_{\rm S}-\sum_{\alpha}\lambda_{\alpha}^{2}\tilde{x}_{\alpha}^{2}-i\sum_{\alpha}\sum_{p}\gamma_{\alpha p}b_{\alpha p}^{+}b_{\alpha p}\\ &+\sum_{\alpha}\sum_{p}\lambda_{\alpha}\left[(\hat{x}_{\alpha}-\tilde{x}_{\alpha})b_{\alpha p}+(\eta_{\alpha p}\hat{x}_{\alpha}-\eta^{*}_{\alpha p}\tilde{x}_{\alpha})b_{\alpha p}^{+}\right].\end{split}$ (8)
@@
To facilitate efficient propagation of the Schrödinger equation in Eq. (5), we employ a tree tensor network state solver.Ke (2023) The high-rank coefficient tensor $C^{\mathbf{n}}_{v_{\rm c}v^{\prime}_{\rm c},\cdots v_{N_{\rm mol}}v^{\prime}_{N_{\rm mol}}}(t)$ is decomposed as a tree tensor network state, which is a powerful way to handle the complicated tensor structure involved in this system. As an example, a graphical representation of the TTNS decomposition of $|\Psi(t)$ for a system with $N_{\rm mol}=8$ molecules and $P=4$ effective environmental modes per bath is shown in Fig. 1b. In this decomposition, a node with $z$-legs corresponds to a rank-$z$ tensor. To achieve an optimal balance between numerical stability, computational efficiency, and simulation time, the number of legs $z$ is constrained to $z\leq 3$.Gunst et al. (2018) The physical indices $v_{\rm c}/v^{\prime}_{\rm c}$ (for the cavity), $v_{i}/v^{\prime}_{i}$ (for the molecules), and $n_{{\rm c}p},\,n_{ip}$ (for the effective environmental modes) in $|\Psi(t)\rangle$ are represented by the dangling legs in the yellow (filled/hollow), red (filled/hollow), blue, and green nodes, respectively. The connected legs between neighboring nodes $j$ and $k$ correspond to virtual indices $r_{jk}$, which runs from $1$ to $D_{jk}$. The maximum value among the bond dimensions $\{D_{jk}\}$ is called the maximal bond dimension $D_{\rm max}$. In practice, $D_{\rm max}$ is systematically increased until the results converge to a desired accuracy. The gray nodes, which have exclusively connected legs, are termed connecting nodes. The root node is assigned to the filled yellow node for the cavity, which has no parent node. All other nodes, except for the root node, have a parent node. This parent node is the adjacent node on the path pointing toward the root. In the binary tree structure displayed in Fig. 1b, each node has at most two children. The leaves are the nodes without any offspring, representing the termination points of the tree.
@@
The super-Hamiltonian $\mathcal{H}$ can be represented as a tree tensor network operator (TTNO) with the same tree topology as the TTNS used for $|\Psi(t)\rangle$. This decomposition ensures that when the TTNO is applied to the TTNS of $\Psi(t)$, the tree structure is preserved. The construction of the TTNO for $\mathcal{H}$ can be performed in an automatic manner.Milbradt, Huang, and Mendl (2024a, b); Ceruti, Kressner, and Sulz (2024); Li et al. (2024) In this work, we use an optimal TTNO constructor based on bipartite graph theory, which allows for an efficient construction of the TTNO.Li et al. (2024) The resulting TTNO, corresponding to the tree shape shown in Fig. 1b, is depicted in Fig. 1d. A key difference between the TTNS and the TTNO is that in the TTNO, the physical nodes possess two dangling legs. To evolve Eq. (5), we adopt a time propagation scheme based on the time-dependent variational principle (TDVP),Haegeman et al. (2016); Paeckel et al. (2019); Dunnett and Chin (2021); Bauernfeind and Aichhorn (2020); Kloss, Reichman, and Bar Lev (2020); Ceruti, Lubich, and Walach (2021) as detailed in Ref. Ke, 2023. After solving Eq. (5), observables of interest can be extracted from the wavefunction $|\Psi(t)\rangle$.
@@
To perform the inner product in Eq. (9), the wavefunction $|{\rm II}\rangle$ is also decomposed as a TTNS, maintaining the same tree topology as $|\Psi(t)\rangle$, as displayed in Fig. 1c. The number next to the legs denotes the bond dimension of the corresponding index. The parameters $d_{\rm m}$, $d_{\rm c}$, and $d_{\rm e}$ represent the number of basis functions for the molecular reactive DoF, cavity mode, and the effective environmental bosons, respectively. Each node in the TTNS is represented by a rank-$3$ tensor $V_{i_{1}\times i_{2}\times i_{3}}$, where $i_{1}$, $i_{2}$, and $i_{3}$ are the dimensions of the respective indices. For the connecting nodes, the tensor is $V_{1\times 1\times 1}=1$. For the molecular (in red) or cavity (in yellow) nodes, the non-zero entries in the tensor are $V_{1\times d_{\rm m/c}\times d_{\rm m/c}}[1,j,k]=\delta_{jk}$ (filled node) and $V_{d_{\rm m/c}\times 1\times d_{\rm m/c}}[j,1,k]=\delta_{jk}$ (hollow node). For the environmental nodes (in green and blue), the non-zero entry is $V_{1\times 1\times d_{\rm e}}[1,1,1]=1$.
@@
In the simulations conducted for this study, the following parameters are used for the double-well model, $E_{b}=2250\,\text{cm}^{-1}$ and $a=44.4$ a.u. , which are consistent with previous studies in the field.Lindoy, Mandal, and Reichman (2023, 2024); Ying and Huo (2023); Hu, Ying, and Huo (2023); Ke, Borrelli, and Thoss (2022); Ke and Richardson (2024b); Fiechter et al. (2023) The Debye-Lorentzian spectral density function is adopted to characterize the baths and their interaction with the system
$J_{\alpha}(\omega)=\frac{2\lambda_{\alpha}^{2}\omega\Omega_{\alpha}}{\omega^{2}+\Omega_{\alpha}^{2}},$ (13)
@@
Figure 2: Ratio of rates $k_{\rm ic}/k_{\rm oc}$ for a single molecule undergoing a reaction inside an optical cavity $(k_{\rm ic})$ compared to the outside rate $(k_{\rm oc})$ as a function of the cavity damping strength $\lambda_{\rm c}$. The results for two different temperatures $T$ and three characteristic frequencies $\Omega_{\rm c}$ of the cavity bath, and light-matter coupling strengths ($\eta_{\rm c}=0.00125$ a.u. in the upper panel and $\eta_{\rm c}=0.005$ a.u. in the lower panel) are shown. The cavity frequency is kept at $\omega_{\rm c}=1185\,\text{cm}^{-1}$ and $d_{\rm c}=10$.
@@ III Results
In practice, the cavity is not perfectly reflective and is therefore subject to external noise from the continuum of far-field electromagnetic modes. Previous numerically exact quantum dynamical studies in the single-molecule limit have demonstrated that, in a resonant but lossless cavity, reaction rates remain unaffected.Lindoy, Mandal, and Reichman (2023); Ying and Huo (2023); Ke and Richardson (2024a) Furthermore, an interpolated rate theory at the single-molecule level, based on Fermi’s Golden Rule in both the lossless and lossy limits, which is valid in the Markovian and weak light-matter coupling regime, predicts a turnover in reaction rates inside the cavity as a function of the cavity loss rate.Ying and Huo (2024) These findings give an important hint that cavity leakage may be a crucial factor in altering reaction dynamics inside an optical microcavity. Accordingly, the objective of this work is to understand how ground-state chemical reactions within an optical cavity are influenced by cavity damping. We begin by investigating the effect of cavity damping strength on reaction rate modifications at the single-molecule level, then extend our analysis to the collective regime, where multiple molecules are simultaneously coupled to the cavity mode.
In this work, cavity leakage is modeled by introducing a cavity bath, with the cavity damping strength quantified by the parameter $\lambda_{\rm c}$. This is verified by examining the dynamics of the average photonic excitation number, given by $\langle n_{\rm c}(t)\rangle=\mathrm{Tr}\{a^{\dagger}_{\rm c}a_{\rm c}\rho(t)\}$. Here, $a^{\dagger}_{\rm c}=\sqrt{\omega_{\rm c}/2}(x_{\rm c}-ip_{\rm c}/\omega_{\rm c})$ and $a_{\rm c}=\sqrt{\omega_{\rm c}/2}(x_{\rm c}+ip_{\rm c}/\omega_{\rm c})$ represent the creation and annihilation operators of the confined cavity mode, respectively. Detailed data supporting this verification are provided in the SI.
@@
Figure 4: Same as Fig. 3 except that $\lambda_{\rm c}=165\,\text{cm}^{-1}$.
@@
Fig. 2 presents the cavity-induced rate modification, represented as $k_{\rm ic}/k_{\rm oc}$–the ratio of reaction rates inside and outside the cavity–in the single-molecule limit, plotted as a function of $\lambda_{\rm c}$. The cavity frequency is set at $\omega_{\rm c}=1185\,\text{cm}^{-1}$, closely resonant with molecular vibrational transitions and corresponding to the peak in the rate modification profile over $\omega_{\rm c}$ (see also Fig. 3b and Fig. 4b). The rate within the cavity initially increases with $\lambda_{\rm c}$, reaching a maximum before this rate enhancement diminishes as $\lambda_{\rm c}$ continues to increase. This behavior is examined with various values of temperature $T$, the characteristic frequency $\Omega_{\rm c}$ of the cavity bath, and the light-matter coupling strength $\eta_{\rm c}$. The extent of the rate modification depends on temperature, $\Omega_{\rm c}$, and $\eta_{\rm c}$. Notably, the cavity-induced rate modification is a finite-temperature effect, vanishing at low temperatures and saturating as the temperature rises, as demonstrated in our previous study.Ke and Richardson (2024a) Furthermore, the rate modification initially increases with $\Omega_{\rm c}$, then plateaus, and gradually decreases at higher $\Omega_{\mathrm{c}}$ values, as demonstrated in the SI. However, both temperature $T$ and $\Omega_{\rm c}$ only slightly affect the crossover value of $\lambda_{\rm c}$. In contrast, increasing the light-matter coupling strength $\eta_{\rm c}$ not only strengthens the rate modification significantly but also shifts the crossover $\lambda_{\rm c}$ to a much higher value, rendering the rate modification more resilient to ambient noise. It is worth noting that, in the strong light-matter coupling regime, the inclusion of higher photonic excitation states is critical. Restricting the consideration to only the photonic ground state and first excited state leads not only to incorrect predictions of the resonant peak in the rate modification profile as a function of the cavity frequency $\omega_{\mathrm{c}}$, but also to a significant overestimation of the turnover $\lambda_{\mathrm{c}}$. The detailed analysis supporting these findings and a more systematic study of the impact of the light-matter coupling strength on reaction dynamics are provided in the SI.
@@
This observation fits into the notion of the stochastic resonance theory, which was first proposed to explain the periodic recurrence of Earth’s ice ages.Benzi, Sutera, and Vulpiani (1981); Nicolis (1981) Over time, it has prospered into a broad field on its own with diverse applications across physics, chemistry, biomedical sciences, engineering, and beyond, highlighting the constructive role of an optimal amount of noise.Wiesenfeld and Moss (1995); Gammaitoni et al. (1998); Andò and Graziani (2000); Hänggi (2002); Harmer, Davis, and Abbott (2002); Wellens, Shatokhin, and Buchleitner (2003); Moss, Ward, and Sannita (2004) Stochastic resonance refers to a phenomenon where a moderate level of noise improves a system’s sensitivity to detect weak, information-carrying signals.Gammaitoni et al. (1998) For chemical reactions inside an optical cavity, three key elements have been possessed for the emergence of stochastic resonance. Firstly, reactions take place in a solvent, which serves as a source of substantial background noise. Secondly, reactive molecules are inherently nonlinear systems that support threshold-like barrier-crossing events. Thirdly, cavity leakage exposes reactions to a generically weak coherent input, which could boost the system’s sensitivity, depending on the noise level.